OpenStudy (anonymous):

find the values of L and M: (6-L)+(3m)i=-12+27i

OpenStudy (anonymous):

for 2 imaginary numbers will equal be equal if and only if both thier real and imaginary parts are equal. \[r1+ i*k1 = r2 +i*k2\] this implies r1=r2, k1=k2